On January 1, 2013, Swifty Corporation issued 2100 of its 10%, $1,000 bonds for $2184000. These bonds were to mature on January 1, 2023 but were callable at 101 any time after December 31, 2016. Interest was payable semiannually on July 1 and January 1. On July 1, 2018, Swifty called all of the bonds and retired them. Bond premium was amortized on a straight-line basis. Before income taxes, Swifty's gain or loss in 2018 on this early extinguishment of debt was
Answer #1

Answer

  • All working forms part of the answer
  • Answer: GAIN in 2018 on early extinguishment of debt = $ 16,800

A

Bonds Face Value

$2,100,000

B

Issue Price

$2,184,000

C = B - A

Premium on Bonds Payable

$84,000

D

Term [years]

10

E = C/20 interest payment

Straight Line amortisation with every interest payment

$4,200

Amortisation of Premium:

01-Jul-13

$4,200

01-Jan-14

$4,200

01-Jul-14

$4,200

01-Jan-15

$4,200

01-Jul-15

$4,200

01-Jan-16

$4,200

01-Jul-16

$4,200

01-Jan-17

$4,200

01-Jul-17

$4,200

01-Jan-18

$4,200

01-Jul-18

$4,200

F

Total Premium Amortised

$46,200

G = C - F

Unamortised Premium when Bonds were called

$37,800

H = A + G

Carrying Value at the time of retirement

$2,137,800

I = A x 101/100

Called at

$2,121,000

J = H - I

Gain on early extinguishment

$16,800 = Answer